OpenStudy (anonymous):

Let f(x) = x2 + 6 and g(x) = x+8/x Find ( g o f)(-7)

OpenStudy (anonymous):

f(x) = x2 + 6 and g(x) = x+8/x f o g is basically putting the g(x) equation into the f(x) equation and x is -7 so f(x) = (x+8/x)^2 + 6 f(-7) = (-7 + 8/-7)^2 +6 = (-7 + -1.14)^2 +6 = (-8.14)^2 +6 = 66.2596 +6 = 72.2596

OpenStudy (anonymous):

ok i get that part. but its not one of my answers

OpenStudy (anonymous):

what are the possible answers? do they have variables in them?

OpenStudy (anonymous):

295/49 384/7 -57/7 63-55

OpenStudy (anonymous):

63/55

OpenStudy (anonymous):

oh. its 63/55 I misread your question. g(x) = x+8/x should look like (x+8)/x basically you plug in -7 into f(x) and then take the answer and plug it in for every x in g(x)

OpenStudy (anonymous):

oh ok